# Exploring the Effect of Explanation Content and Format on User Comprehension and Trust

### v1

@@ 5. Discussion and Limitations
Across two studies with different samples (general population and medical students), we found a reliable preference for occlusion over Shapley value explanations. Our participants reported that they understood and trusted the occlusion explanations more than the SHAP explanations, as we expected. However, the comparison of all the explanations revealed that this effect may be solely driven by the fact that our participants found the OT explanations consistently more comprehensible than the SC explanations. Moreover, in Study 1, the OT explanations were also preferred to the SC explanations, though in Study 2 the results were less clear as there was no evidence that the former explanations were preferred over the latter. These findings are consistent with a notion of people preferring explanations in text over chart format (Bertrand et al., 2023), for instance, due to finding some graphs difficult to interpret (Durand et al., 2020; Galesic and Garcia-Retamero, 2011; Szymanski et al., 2021). This idea somewhat undermines the potential message that, in terms of content, FA explanations based on simple rather than complex models should be prioritised. Interestingly, the results regarding trustworthiness were slightly in conflict with this idea, since in our medical student sample we only found evidence for higher trust in the OC compared to the SC explanations. While these findings show some support for the role of explanation content, more research is needed to explore how people relate to and utilise different XAI techniques, and how and when an explanation’s format (e.g. text vs. chart) trumps the impact of its content. Future research should also explore the potential of tailoring explanation content and format to individual needs (Bhatt et al., 2020), such as their expertise in a field (Wang and Yin, 2023) or AI-systems (Schoeffer et al., 2022) and their skills to interpret numerical information (e.g. graph literacy (Galesic and Garcia-Retamero, 2011; Durand et al., 2020) or numeracy (Reyna et al., 2009; Gleaves et al., 2020)).
Our novel performance measure (i.e. the definition recognition task) could also not settle the question regarding the impact of content as we found no evidence that the participants were better at recognising the definition of either explanation. Nonetheless, the lack of positive findings could be related to how we measured performance. That is, the task of identifying the correct definition of explanations may have been too difficult as performance was only reliably over chance level for the occlusion explanations in the medical student sample potentially creating a floor effect for our measure. Future work should operationalise performance in other formats also.
Our studies carry some limitations. Both studies lacked a SHAP explanation in text format and so a complete design, making it difficult to fully separate the impact of the content of an explanation from the format. Nonetheless, as argued previously, it is not obvious how an ST explanations that is fairly represenative of SHAP can be created. Moreover, the relatively lower sample size of Study 2 may have caused some of our tests to be insensitive and non-significant preventing us from settling the question of whether occlusion explanations are perceived to be more comprehensible than SHAP explanations regardless of the format. For example, a significant difference between OC and SC explanations in terms of their comprehensibility would challenge the idea that preference for occlusion explanations is driven by a preference for charts. Finally, self-report measures are prone to biases, and a recent study (Nagendran et al., 2023) found no correlation between self-reported understanding of XAI and the impact of XAI on behaviour, casting some doubts on their generalisability to real-life behaviour. Future research should investigate this presumed dissociation and focus on behavioural measures of advice uptake (Pálfi et al., 2022; Pálfi et al., 2023; Panigutti et al., 2022) and understanding of advice.

### v2

@@ 5 Discussion
Across two studies with different samples (from the general population and medical students), we found a reliable preference for Occlusion-1 over SHAP explanations. Our participants reported that they understood and trusted the former more than the latter, as expected. However, the comparison of all the explanations revealed that this effect may be solely driven by the fact that our participants found the OT explanations consistently more comprehensible than the SC explanations. Moreover, in Study 1, the OT explanations were also preferred to the SC explanations, though in Study 2 the results were less clear. These findings are consistent with a notion of people, or more specifically patients within a healthcare setting, preferring explanations in text over chart format Bertrand et al. (2023), for instance, due to finding some graphs difficult to interpret Durand et al. (2020); Szymanski et al. (2021). These findings indicate that prioritising FA explanations based on simple rather than complex models may not contribute to clarity. Interestingly, the results regarding trustworthiness were slightly in conflict with this idea, since in our medical student sample we only found evidence for higher trust in the OC compared to the SC explanations. While these findings show support for the role of explanation content, more research is needed to explore how people relate to and utilise different XAI techniques, and how and when an explanation’s format trumps the impact of its content. Future research should also explore the potential of tailoring explanation content and format to individuals Lesley and Hernández (2024), e.g. based on their expertise in a field Wang and Yin (2023) and their skills in interpreting numerical information (e.g. graph literacy Durand et al. (2020) or numeracy Reyna et al. (2009)).
Our studies did not assess SHAP explanations in text format, since, as we discussed, it is not obvious how to create an “ST” explanation that is representative of SHAP and comprehensible. In any case, even if we could have included these additional explanations, the relatively lower sample size of Study 2 may have caused some of our tests to be insensitive and non-significant, preventing us from settling the question of whether Occlusion-1 explanations are perceived to be more comprehensible than SHAP explanations regardless of the format, e.g. challenging the idea that the preference for Occlusion-1 explanations is driven by a preference for charts.
Our expert sample in Study 2 comprised medical students so it is not obvious how well our findings generalise to experienced healthcare practitioners: we leave this to future work. Also, self-report measures are prone to biases: Nagendran et al. Nagendran et al. (2023) found no correlation between self-reported understanding of XAI and the impact of XAI on behaviour, casting some doubt on their generalisability to real-life behaviour. Future research should investigate this presumed dissociation for behavioural measures of advice understanding and uptake Pálfi et al. (2022); Panigutti et al. (2022). This justifies our definition of a novel performance measure (i.e. the definition recognition task), especially since we found no evidence that participants were better at recognising the definition of either explanation: it would be interesting to study variants of this measure or different ways to assess it to better settle the question of the impact of explanation content. The task of choosing the correct definition of explanations may have been too difficult as performance was only reliably over chance level for the Occlusion-1 explanations in the medical student sample, potentially creating a floor effect, but, given the simple descriptions to choose from, this in itself calls into question user comprehension of FA explanations more generally.
In summary, by evaluating explanations across various realistic scenarios, metrics, and user groups, our approach makes a novel contribution to the understanding of how XAI techniques can be best used in healthcare. While our experiments did not provide definitive evidence for the dominance of Occlusion-1 over SHAP explanations, this lack of a clear preference is interesting in itself, especially as the simpler Occlusion-1 explanations were expected to be favoured. These findings underscore the importance of further research and empirical evaluation on the interplay between explanation content and format, pointing to several future directions.
Ethical Statement
Participation in both studies was voluntary. Participants were asked to provide informed consent after reading an information sheet and they could withdraw at any time prior to completion, simply by closing the survey window, with their data deleted permanently. We informed the participants that once they had completed the survey, they could no longer withdraw and we would keep their data. All the participant responses were anonymised once payment was made (i.e. we deleted their e-mail addresses from the dataset). We received ethics approval from our institution for these studies.
@@
Gender Age BMI SS T2 CP LA WL DS AP BV I H A BH
Ex. 2 M 60 37.55 light yes no no yes no yes no no no no no
Ex. 3 F 74 34.60 ex no no yes no yes no yes no yes no no
Ex. 4 F 60 27.85 non no no yes yes yes no no no no no no
Table 4: Values used for patients described in Examples 2-4 (See Table 1 for the variables).